A: The activation energy of a certain reaction is 36.8 kJ/mol . At 27  ∘C , the rate constant is 0.0120s−1. At what temperature in degrees Celsius would this reaction go twice as fast?

B: Given that the initial rate constant is 0.0120s−1 at an initial temperature of 27  ∘C , what would the rate constant be at a temperature of 100.  ∘C for the same reaction described in Part A?
Answer #1

since we know that by the arehenius equation

ln(K2/K1) = Ea/R*(1/T1- 1/T2)

K2 = 2xK1

ln(2) = 36800/8.314 * ( 1/(27+273) - 1/(T+273))

T = -( ln(2) / (36800) * 8.314 - 1/(27+273))^-1

T = 314.8 - 273 = 41.8 C

b)

ln(k2/(0.012)) = 36800/8.314 * ( 1/(27+273) - 1/(100+273))

k2 = 0.012*exp( 36800/8.314 * ( 1/(27+273) - 1/(100+273)))

K2 = 0.215 1/s